Please help, I'm not good at word problems.
Two of the top grossing concert tours were by a jazz band and a rock band. Together, the two tours visited 174 cities. The jazz band visited 96 cities more than the rock band.
How many cities did each group visit.
Jazz Band
Rock Band

Found 2 solutions by checkley77, mastermath:
Answer by checkley77(12844)

x+(x+96)=174
2x=174-96
2x=78
x=78/2
x=39 cities visited by the rock band.
39+96=135 cities visited by the jazz band.
Proof
39+135=174
174=174

Answer by mastermath(14)

Let the number of cities visited by Rock band be x.
The jazz band visited 96 cities more than the rock band.
So, the number of cities visited by Jazz band will be (x+96).
Together, the two tours visited 174 cities.
So, add both of them
Rock + Jazz = 174
x+(x+96)=174
2x+96=174
subtract 96 on both the sides
2x+96-96=174-96
2x=78
Divide with 2 on both the sides
x=39
so, x+96 = 39+96 = 135
Thus,the number of cities visited by Rock band is 39.
And the number of cities visited by Jazz band is 135.